# Exploiting Multimodal Synthetic Data for Egocentric Human-Object Interaction Detection in an Industrial Scenario

### v1

@@ Abstract
In this paper, we tackle the problem of Egocentric Human-Object Interaction (EHOI) detection in an industrial setting. To overcome the lack of public datasets in this context, we propose a pipeline and a tool for generating synthetic images of EHOIs paired with several annotations and data signals (e.g., depth maps or instance segmentation masks). Using the proposed pipeline, we present EgoISM-HOI a new multimodal dataset composed of synthetic EHOI images in an industrial environment with rich annotations of hands and objects. To demonstrate the utility and effectiveness of synthetic EHOI data produced by the proposed tool, we designed a new method that predicts and combines different multimodal signals to detect EHOIs in RGB images. Our study shows that exploiting synthetic data to pre-train the proposed method significantly improves performance when tested on real-world data. Moreover, the proposed approach outperforms state-of-the-art class-agnostic methods. To support research in this field, we publicly release the datasets, source code, and pre-trained models at https://iplab.dmi.unict.it/egoism-hoi.
@@ 2 Related Work
2.1 Datasets for Human-Object Interaction Detection from Third Person View
Previous works have proposed benchmark datasets to study human-object interactions from a third-person vision. PASCAL VOC (Everingham et al., 2009) was one of the first datasets focusing on understanding human behavior from images. This dataset has been used for several tasks related to human-object interaction understanding, such as object classification, object detection, and static action classification. Gupta and Malik (2015) introduced The Verbs in COCO (V-COCO) dataset, an extension of the COCO dataset (Lin et al., 2014) that includes 26 verbs classes, along with bounding box annotations of human and objects involved in interactions. Chao et al. (2015) presented Humans Interacting with COmmon Objects (HICO), a dataset for detecting human-object interactions that comprises more than 600 categories of human-object interactions across 117 activities and 80 common objects. Chao et al. (2018) extended HICO to HICO-DET, adding more than 150,000 annotated instances of human-object interaction pairs. Li et al. (2020) proposed AmbiguousHOI, a benchmark that includes hard ambiguous images of HOI instances selected from existing datasets such as HICO-DET (Chao et al., 2018), V-COCO (Gupta and Malik, 2015), and OpenImage (Kuznetsova et al., 2020). The Human-Object Interaction for Application (HOI-A) dataset has been proposed by Liao et al. (2020). It includes 38,668 annotated images with 11 different types of objects, and 10 action categories, comprising 43,820 human instances, 60,438 object instances, and 96,160 interaction instances. Recently, the human-object interaction detection task has been studied by exploiting multimodal signals. The BEHAVE dataset (Bhatnagar et al., 2022) is a multi-view RGB-D dataset of human-object interactions acquired in natural environments, which contains 3D human and object annotations, instance segmentation masks, and contact-state labels. Most related to our study, 100 Days of Hands (Shan et al., 2020) is a large-scale dataset of human-object interactions containing more than 131 days of video footage acquired from both third and first-person points of view. The authors extracted 100K frames and annotated with bounding boxes 189.6K hands and 110.1K objects involved in interactions. Moreover, for each hand, they annotated the contact state considering five different classes (i.e., none, self, other-person, non-portable object, and portable object).
The aforementioned works focused mostly on a third-person point of view. Our study focuses on understanding human-object interactions from a first-person point of view.
2.2 Datasets for Human-Object Interaction Detection from First Person View
@@ 2.4 Methods for Detecting Human-Object Interactions
In the past year, the human-object interaction detection task has been studied from the third-person point of view (Gupta and Malik, 2015; Chao et al., 2015, 2018). Gkioxari et al. (2018) proposed a method for detecting human-object interactions in the form of $<$human, verb, object$>$ triplets, where bounding boxes around objects and humans are also predicted. Specifically, they extended the state-of-the-art object detector Faster R-CNN (Ren et al., 2015) with an additional human-centric branch that uses the features extracted by the backbone to predict a score for candidate human-object pairs and an action class. Liao et al. (2020) proposed a method called PPDM (Parallel Point Detection and Matching) that defines an HOI as a triplet $<$human point, interaction point, object point$>$ composed of three points associated with the human, the active object and the interaction location. Recently, several works figured out the HOI detection task by proposing transformer-based models. Zhang et al. (2022a) proposed a new two-stage detector based on a transformer architecture to detect interactions. Wu et al. (2022) proposed an approach for learning a body-part saliency map, which contains informative cues of the person involved in the interaction and other persons in the image, in order to boost HOI detection methods (Chao et al., 2018; Gao et al., 2018). Ma et al. (2023) introduced a transformer-based human-object interaction detector that uses a multi-scale feature extractor and a multi-scale sampling strategy to predict the HOI instances from images with noisy backgrounds in the form of $<b_{h},b_{o},c_{o},c_{v}>$ quadruplet, where $b_{h}$ and $b_{o}$ represent the human and object boxes, and $c_{o}$ and $c_{v}$ the object class and the verb class. While previous works all addressed the HOI modeling detecting a bounding box around the human, Shan et al. (2020) addressed the HOI detection task by predicting information about human hands, such as hand location, side, contact state, and, in case of an interaction, a box around the object touched by the hand. Zhang et al. (2022b) proposed to use a contact boundary, i.e. the contact region between the hand and the interacting object, to model the interaction relationship between hands and objects. Fu et al. (2022) designed an approach for HOI detection that introduced a new pixel-wise voting function for improving the active object bounding box estimation. Benavent-Lledo et al. (2022) proposed an architecture for human-object interaction detection estimation based on two YOLOv4 object detectors (Bochkovskiy et al., 2020) and an attention-based method. Recently, some work investigated the use of additional modalities, such as 6DOF hand poses or semantic segmentation masks, to learn more robust representations of human-object interactions. Lu and Mayol-Cuevas (2021) introduced an approach that uses contextual information, i.e. hand pose, hand mask, and object mask, to improve the performance of HOI detection systems.
@@ 3 Proposed EHOI Generation Pipeline
We developed a tool based on the Unity [^4] https://unity.com/ engine which exploits 3D models of the objects and the environment to generate synthetic egocentric human-object interaction images together with the following data: 1) RGB images (see Fig. 4 - left), 2) depth maps (see Fig. 4 - right), 3) instance segmentation masks (see Fig. 4 - center), 4) bounding boxes for hands and objects including the object categories, 5) EHOI’s metadata, such as information about associations between hands and objects in contact (which hand is in contact with which object), and hand attributes (i.e., hand side, and hand contact state).
@@ 4 EgoISM-HOI dataset
For EgoISM-HOI-Real, we collected and labeled 42 real egocentric videos in the ENIGMA Laboratory. In these videos, subjects performed testing and repairing operations on electrical boards using laboratory tools. To simplify data collection and to make it more consistent, we developed an application for Microsoft Hololens 2666https://www.microsoft.com/hololens that guides the users through audio and images, suggesting the next steps to perform during the acquisition. We defined 8 procedures composed of several steps, in which we vary the tools and electrical boards interacted by the users. Nineteen subjects participated in the data collection. Two were women and seventeen were men. For privacy reasons, we made sure that no other people are visible in the videos, and all the subjects removed any personal object that might reveal their identities (e.g., rings or wristwatches). We acquired 18 hours, 48 minutes, and 13 seconds of video recordings, with an average duration of 26 minutes and 51 seconds, at a resolution of 2272x1278 pixels and a framerate of 30fps. Table 2 summarizes statistics about the collected data.
@@ 5 Proposed approach
Inspired by Shan et al. (2020), our system extends a two-stage object detector with additional modules specialized to recognize human-object interactions. The proposed method is able to exploit different data signals, such as instance segmentation maps and depth maps, to improve the performance of classic HOI detection approaches. Moreover, our method is able to recognize the class of all the objects in the scene. We believe that this knowledge could be used for other downstream tasks.
@@
Table 3: Results of the proposed approach on EgoISM-HOI-Real test data. The EgoISM-HOI-Synth column indicates whether the EgoISM-HOI-Synth training set was used for pre-training models. The EgoISM-HOI-Real % column shows the percentage of real-world data used for fine-tuning.
EgoISM-HOI-Synth EgoISM-HOI-Real % AP Hand AP H.+Side AP H.+State mAP H.+Obj mAP H.+All
@@
No 10 90.08 88.57 45.69 18.19 17.48
No 25 90.43 89.45 43.73 18.72 18.31
No 50 90.43 89.57 52.74 19.17 19.06
No 100 90.54 90.06 56.34 22.31 21.76
@@ 6 Experimental results
We conducted a series of experiments to 1) assess how much the generated synthetic data are useful in training models able to generalize to the real-world domain (Section 6.2), 2) highlight the contribution of multimodal signals to tackle the EHOI detection task (Section 6.3), and 3) compare the proposed method with a set of baselines based of state-of-the-art class-agnostic approaches (Section 6.4). Section 6.5 further reports additional results on pre-training our method with external data and improvements obtained by our approach on the object detection task.
@@ 6.2 The Impact of Synthetic Data on System Performance
Table 3 reports EHOI detection results on the EgoISM-HOI-Real test set. Models pre-trained with EgoISM-HOI-Synth data (rows 1-5) outperform all the corresponding models trained using only EgoISM-HOI-Real data (rows 6-9) by consistent margins according to all evaluation metrics, except for the AP Hand+Side measure, in which they obtain the second-best result (row 4). Considering the two models fine-tuned using the 100% of the real-world training set (rows 5 and 9), the improvements of the model pre-trained with EgoISM-HOI-Synth data are significant in the metrics affected by active objects, obtaining +13,63% (35.94 vs 22.31) for the mAP Hand+Obj and +13,71% (35.47 vs 21.76) for the mAP Hand+All. These improvements are also evident if we compare the models trained with smaller portions of the real-world data. However, for the metrics AP Hand and AP Hand+State, there is only a small boost in the performance of the model pre-trained on EgoISM-HOI-Synth (row 5) compared to the model trained using only EgoISM-HOI-Real data (row 9), i.e., +0.19% (90.73 vs 90.54) and +0.54% (56.88 vs 56.34). These results suggest that using synthetic data for pre-training models enhances the method’s ability to detect active objects which are susceptible to frequent occlusions by the hands. In addition, it is worth noting that the model trained using only the EgoISM-HOI-Synth data (row 1) outperforms the best model that used only the real-world data for the evaluation measures influenced by the active objects, obtaining +1,61% (23.92 vs 22.31) and +1,52% (23.28 vs 21.76) for the mAP Hand+Obj and mAP Hand+All metrics, respectively. Figure 9 further illustrates the results in terms of mAP Hand+All considering different amounts of EgoISM-HOI-Real training data in the fine-tuning.
@@ 6.3 Impact of Multimodal training
6.4 Comparison with class-agnostic baselines
Table 5 compares our system with different instances of the class-agnostic method introduced in Shan et al. (2020). Henceforth, we will refer to this method as Hands In Contact (HIC). Since HIC is class agnostic, to compare our method with it, we extend it to recognize the active object classes following two different approaches. In the first approach, we used a Resnet-18 CNN (He et al., 2016b) to classify image patches extracted from the active object bounding boxes. We trained the classifier with four different sets of data: 1) BS1: we sampled 20,000 frames from 19 videos where a single object of each class is shot at a time. This collection provides a minimal training set that can be collected with a modest labeling effort (comparable with the time needed for acquiring 3D models of the objects in our pipeline); 2) BS2: we used images from the proposed EgoISM-HOI-Real training set; 3) BS3: we used images from the proposed EgoISM-HOI-Synth training set; 4) BS4: we used all EgoISM-HOI data. The second approach (BS5) exploits a YOLOv5999YOLOv5: https://github.com/ultralytics/yolov5 object detector, trained to recognize the considered objects (see Fig. 3), to assign a label to the active objects predicted by HIC. Specifically, for each active object prediction, we select the class of the object with the highest IoU among those predicted by the YOLOv5 object detector or discard the proposal if there are no box intersections. It is worth noting that HIC was pre-trained on the large-scale dataset 100DOH, which contains over 100K labeled frames of HOIs.
@@ 6.4 Comparison with class-agnostic baselines
Method EgoISM-HOI-Synth EgoISM-HOI-Real % mAP Hand+All
@@
The best model of the proposed EHOI detection method (row 3) outperforms all the baselines (rows 4-8) with significant improvements ranging from +12.92% (36.51 vs 23.59) to +26,59% (36.51 vs 9.92). The approach based on Resnet-18 (rows 4-7) leads to better performance compared to the method based on the YOLOv5 object detector (row 8). Indeed, considering only the baselines (rows 4-8), the best result is achieved by BS4 (row 7), which was pre-trained using synthetic and real-world EgoISM-HOI data, with an improvement of +2.97% (23.59 vs 20.62) over the BS5 (row 8). Interestingly, even the BS2 (row 5), which did not use synthetic data during training, obtained a higher result of +1.56% (22.18 vs 20.62) than the BS5 (row 8). These results suggest the limits of this simple approach. In addition, it is worth noting that the model pre-trained on EgoISM-HOI-Synth and fine-tuned using 10% of the EgoISM-HOI-Real training set (row 2) outperforms all the baseline approaches (rows 4-8), with an improvement of +7,06% (30.65 vs 23.59) over the BS4 (row 7). Worth mentioning that the model trained only on EgoISM-HOI-Synth (row 1) achieves comparable results to the best baseline approach (row 7).
6.5 Additional results
In this section, we show an additional set of experiments with the aim of 1) demonstrating how using domain-specific synthetic data improves the performance of a system pre-trained on an out-of-domain large-scale dataset (Section 6.5.1), 2) showing the potential of using synthetic data for the related task of Object Detection (Section 6.5.2). Similar to the set of experiments in Section 6.2, we deactivated the multimodal hand state classifier, monocular depth estimation branch, and instance segmentation branch modules for a fair comparison.
6.5.1 Pre-training on 100 Days Of Hands
To further demonstrate the utility of synthetic data, we performed an additional experiment in which we pre-trained different models on the large-scale 100DOH dataset and then fine-tuned them on our EgoISM-HOI dataset. The goal of this experiment is to demonstrate how the use of domain-specific synthetic data further increases the performance of a system pre-trained on a large amount of out-of-domain real data.

### v2

@@ Abstract
In this paper, we tackle the problem of Egocentric Human-Object Interaction (EHOI) detection in an industrial setting. To overcome the lack of public datasets in this context, we propose a pipeline and a tool for generating synthetic images of EHOIs paired with several annotations and data signals (e.g., depth maps or segmentation masks). Using the proposed pipeline, we present EgoISM-HOI a new multimodal dataset composed of synthetic EHOI images in an industrial environment with rich annotations of hands and objects. To demonstrate the utility and effectiveness of synthetic EHOI data produced by the proposed tool, we designed a new method that predicts and combines different multimodal signals to detect EHOIs in RGB images. Our study shows that exploiting synthetic data to pre-train the proposed method significantly improves performance when tested on real-world data. Moreover, to fully understand the usefulness of our method, we conducted an in-depth analysis in which we compared and highlighted the superiority of the proposed approach over different state-of-the-art class-agnostic methods. To support research in this field, we publicly release the datasets, source code, and pre-trained models at https://iplab.dmi.unict.it/egoism-hoi.
@@ 2 Related Work
2.1 Datasets for Human-Object Interaction Detection
Previous works have proposed benchmark datasets to study human-object interactions from a third-person vision. The datasets, such as PASCAL VOC (Everingham et al., 2009), V-COCO (Gupta and Malik, 2015), HICO (Chao et al., 2015), HICO-DET (Chao et al., 2018), AmbiguousHOI (Li et al., 2020), HOI-A (Liao et al., 2020), and BEHAVE (Bhatnagar et al., 2022), offer diverse annotations and cover a wide range of scenarios. Most related to our study is 100 Days of Hands (Shan et al., 2020) which is a large-scale dataset of human-object interactions containing more than 131 days of video footage acquired from both third and first-person points of view. The authors extracted 100K frames and annotated with bounding boxes 189.6K hands and 110.1K objects involved in interactions. Moreover, for each hand, they annotated the contact state considering five different classes (i.e., none, self, other-person, non-portable object, and portable object). Differently from previous works, our study focuses on understanding human-object interactions from a first-person point of view with the exploitation of synthetic generated data.
@@ 2.3 Methods for Detecting Human-Object Interactions
In the past years, the human-object interaction detection task has been studied from the third-person point of view (Gupta and Malik, 2015; Chao et al., 2015, 2018). Gkioxari et al. (2018) proposed a method for detecting human-object interactions in the form of $<$human, verb, object$>$ triplets, where bounding boxes around objects and humans are also predicted. Specifically, they extended the state-of-the-art object detector Faster R-CNN (Ren et al., 2015) with an additional human-centric branch that uses the features extracted by the backbone to predict a score for candidate human-object pairs and an action class. Liao et al. (2020) proposed a method called PPDM (Parallel Point Detection and Matching) that defines an HOI as a triplet $<$human point, interaction point, object point$>$ composed of three points associated with the human, the active object and the interaction location. Recently, several works figured out the HOI detection task by proposing transformer-based models. Zhang et al. (2022a) proposed a new two-stage detector based on a transformer architecture to detect interactions. Wu et al. (2022) proposed an approach for learning a body-part saliency map, which contains informative cues of the person involved in the interaction and other persons in the image, in order to boost HOI detection methods (Chao et al., 2018; Gao et al., 2018). Ma et al. (2023) introduced a transformer-based human-object interaction detector that uses a multi-scale feature extractor and a multi-scale sampling strategy to predict the HOI instances from images with noisy backgrounds in the form of $<b_{h},b_{o},c_{o},c_{v}>$ quadruplet, where $b_{h}$ and $b_{o}$ represent the human and object boxes, and $c_{o}$ and $c_{v}$ the object class and the verb class. While previous works all addressed the HOI modeling detecting a bounding box around the human, Shan et al. (2020) addressed the HOI detection task by predicting information about human hands, such as hand location, side, contact state, and, in case of an interaction, a box around the object touched by the hand. Zhang et al. (2022b) proposed to use a contact boundary, i.e. the contact region between the hand and the interacting object, to model the interaction relationship between hands and objects. Fu et al. (2022) designed an approach for HOI detection that introduced a new pixel-wise voting function for improving the active object bounding box estimation. Benavent-Lledo et al. (2022) proposed an architecture for human-object interaction detection estimation based on two YOLOv4 object detectors (Bochkovskiy et al., 2020) and an attention-based method. Recently, some work investigated the use of additional modalities, such as 6DOF hand poses or semantic segmentation masks, to learn more robust representations of human-object interactions. Lu and Mayol-Cuevas (2021) introduced an approach that uses contextual information, i.e. hand pose, hand mask, and object mask, to improve the performance of HOI detection systems.
@@ 3 Proposed EHOI Generation Pipeline
We developed a tool based on the Unity [^4] https://unity.com/ engine which exploits 3D models of the objects and the environment to generate synthetic egocentric human-object interaction images together with the following data: 1) RGB images (see Fig. 4 - left), 2) depth maps (see Fig. 4 - right), 3) instance segmentation masks (see Fig. 4 - center), 4) bounding boxes for hands and objects including the object categories, 5) EHOI’s metadata, such as information about associations between hands and objects in contact (which hand is in contact with which object), and hand attributes (i.e., hand side, and hand contact state). Differently from our previous work (Leonardi et al., 2022), the new tool streamlines the setup of virtual scenes, enhances interaction realism, facilitates the generation of diverse modalities, introduces support for RGB video generation, and allows customization of various scene parameters.
@@ 4 EgoISM-HOI dataset
For EgoISM-HOI-Real, we expand the previous original real-world dataset (Leonardi et al., 2022) from 8 to 42 real egocentric videos in the ENIGMA Laboratory. In these videos, subjects performed testing and repairing operations on electrical boards using laboratory tools. We developed an application for Microsoft Hololens 2, designed to assist operators through the data acquisition process. This application offers audio guidance and shows images to facilitate complex operations during the acquisition phase. Additionally, it integrates voice commands to enhance human-device interaction, allowing operators to give commands such as ”next” or ”back” to navigate through procedure steps. We defined 8 procedures composed of several steps, in which we vary the tools and electrical boards interacted by the users. Nineteen subjects participated in the data collection. Two were women and seventeen were men. For privacy reasons, we made sure that no other people were visible in the videos, and all the subjects removed any personal object that might reveal their identities (e.g., rings or wristwatches). We acquired 18 hours, 48 minutes, and 13 seconds of video recordings, with an average duration of 26 minutes and 51 seconds, at a resolution of 2272x1278 pixels and a framerate of 30fps. Table 2 summarizes statistics about the collected data.
@@ 5 Proposed approach
Inspired by Shan et al. (2020), our system extends a two-stage object detector with additional modules specialized to recognize human-object interactions. Differently from our previous work (Leonardi et al., 2022), the proposed method is able to exploit different data signals, such as instance segmentation maps and depth maps, to improve the performance of classic HOI detection approaches. Moreover, our method is able to recognize the class of all the objects in the scene. We believe that this knowledge could be used for other downstream tasks.
@@
Table 3: Results of the proposed approach on EgoISM-HOI-Real test data. The EgoISM-HOI-Synth column indicates whether the EgoISM-HOI-Synth training set was used for pre-training models. The EgoISM-HOI-Real % column shows the percentage of real-world data used for fine-tuning. The improvement rows show the improvements of models pre-trained with synthetic data compared to models using only real data.
EgoISM-HOI-Synth EgoISM-HOI-RealAP Hand AP H.+Side AP H.+State mAP H.+Obj mAP H.+All
@@
Improvement over 100% EgoISM-HOI-Real only +0.19 ${\color[rgb]{1,0,0}-0.07}$ +0.54 +13.63 +13.71
@@ 6 Experimental results
We conducted a series of experiments to 1) assess how much the generated synthetic data are useful in training models able to generalize to the real-world domain (Section 6.2), 2) highlight the contribution of multimodal signals to tackle the EHOI detection task (Section 6.3), and 3) compare the proposed method with a set of baselines based on state-of-the-art class-agnostic approaches (Section 6.4). Section 6.5 reports ablation studies, delving into the contributions of various modules and the impact of different volumes of synthetic images on our model’s performance. Moreover, it reports additional results on pre-training our method with external data and improvements obtained by our approach for the object detection task.
@@ 6.2 The Impact of Synthetic Data on System Performance
Table 3 reports EHOI detection results on the EgoISM-HOI-Real test set. Models pre-trained with EgoISM-HOI-Synth data outperform all the corresponding models trained using only EgoISM-HOI-Real data by consistent margins according to all the evaluation metrics. Considering the two models fine-tuned using the 100% of the EgoISM-HOI-Real training set, the improvements of the model pre-trained with EgoISM-HOI-Synth data are significant in the evaluation measures affected by active objects (i.e., mAP Hand+Obj and mAP Hand+All). Specifically, there is an increase of +13.63% (35.94% vs. 22.31%) for mAP Hand+Obj and of +13.71% (35.47% vs. 21.76%) for mAP Hand+All. The improvements persist across all the other configurations of real-world training data, i.e., 10%, 25%, and 50%. Models pre-trained with synthetic data show considerable increments of performances of +13.17%, +13.02%, and +14.06%, respectively, in terms of mAP Hand+All, compared to models trained only on EgoISM-HOI-Real data.
Considering the AP Hand, AP H.+Side and AP Hand+State evaluation measures, we observe marginal enhancements in the performance of models pre-trained on EgoISM-HOI-Synth. These results suggest that using synthetic data for pre-training models significantly improves the method’s capability to detect active objects, which are susceptible to frequent occlusions by the hands. Measures influenced only by hands show minimal benefit from additional real-world data, suggesting that they reach a saturation point earlier in terms of performance improvement.
In addition, it is worth noting that the model trained using only the EgoISM-HOI-Synth data (row 1) outperforms the best model that used only the real-world data for the evaluation measures influenced by the active objects, obtaining +1,61% (23.92% vs 22.31%) and +1,52% (23.28% vs 21.76%) for the mAP Hand+Obj and mAP Hand+All measures respectively. Figure 9 further illustrates the results in terms of mAP Hand+All considering different amounts of EgoISM-HOI-Real training data in the fine-tuning.
@@ 6.3 Impact of Multimodal training
Method EgoISM-HOI-Synth EgoISM-HOI-RealmAP Hand+All
@@
6.4 Comparison with class-agnostic baselines
This section compares our proposed approach with different baseline approaches based on state-of-the-art methods (Shan et al., 2020; Darkhalil et al., 2022). Specifically, we compare our approach with the method of Shan et al. (2020) in section 6.4.1 and with the method of Darkhalil et al. (2022) in section 6.4.2.
6.4.1 Comparison with HiC
Table 5 compares our system with different instances of the class-agnostic method introduced in Shan et al. (2020). Henceforth, we will refer to this method as Hands In Contact (HIC). Since HIC is class agnostic, to compare our method with it, we extend it to recognize the active object classes following two different approaches. In the first approach, we used a Resnet-18 CNN (He et al., 2016b) to classify image patches extracted from the active object bounding boxes. We trained the classifier with four different sets of data: 1) BS1: we sampled 20,000 frames from 19 videos where a single object of each class is shot at a time. This collection provides a minimal training set that can be collected with a modest labeling effort (comparable with the time needed for acquiring 3D models of the objects in our pipeline); 2) BS2: we used images from the proposed EgoISM-HOI-Real training set; 3) BS3: we used images from the proposed EgoISM-HOI-Synth training set; 4) BS4: we used all EgoISM-HOI data. The second approach (BS5) exploits a YOLOv5888YOLOv5: https://github.com/ultralytics/yolov5 object detector, trained to recognize the considered objects (see Fig. 3), to assign a label to the active objects predicted by HIC. Specifically, for each active object prediction, we select the class of the object with the highest IoU among those predicted by the YOLOv5 object detector or discard the proposal if there are no box intersections. It is worth noting that HIC was pre-trained on the large-scale dataset 100DOH, which contains over 100K labeled frames of HOIs.
@@ 6.4.1 Comparison with HiC
The best model of the proposed EHOI detection method (row 3) outperforms all the baselines (rows 4-8) with significant improvements ranging from +12.92% (36.51 vs 23.59) to +26,59% (36.51 vs 9.92). The approach based on Resnet-18 (rows 4-7) leads to better performance compared to the method based on the YOLOv5 object detector (row 8). Indeed, considering only the baselines (rows 4-8), the best result is achieved by BS4 (row 7), which was pre-trained using synthetic and real-world EgoISM-HOI data, with an improvement of +2.97% (23.59 vs 20.62) over the BS5 (row 8). Interestingly, even the BS2 (row 5), which did not use synthetic data during training, obtained a higher result of +1.56% (22.18 vs 20.62) than the BS5 (row 8). These results suggest the limits of this simple approach. In addition, it is worth noting that the model pre-trained on EgoISM-HOI-Synth and fine-tuned using 10% of the EgoISM-HOI-Real training set (row 2) outperforms all the baseline approaches (rows 4-8), with an improvement of +7,06% (30.65 vs 23.59) over the BS4 (row 7). It is worth mentioning that the model trained only on EgoISM-HOI-Synth (row 1) achieves comparable results to the best baseline approach (row 7).
6.4.2 Comparison with VISOR baseline
Table 6: Comparison between the proposed approach and VISOR baseline (Shan et al., 2020) on EgoISM-HOI and VISOR datasets.
Method Training set Test set AP Hand+All
VISOR baseline EgoISM-HOI-Synth EgoISM-HOI-Real test 10.79
Proposed EgoISM-HOI-Synth EgoISM-HOI-Real test 20.47
Improvement +9.68
VISOR baseline EgoISM-HOI-Real EgoISM-HOI-Real test 26.58
Proposed EgoISM-HOI-Real EgoISM-HOI-Real test 33.17
Improvement +6.59
VISOR baseline EgoISM-HOI EgoISM-HOI-Real test 27.60
Proposed EgoISM-HOI EgoISM-HOI-Real test 32.13
Improvement +4.53
VISOR baseline VISOR VISOR validation 44.76
Proposed VISOR VISOR validation 46.97
Improvement +2.21
Table 6 compares our framework with respect to the class-agnostic HOI detection baseline introduced in Darkhalil et al. (2022). To comprehensively assess the effectiveness of our proposed approach, we conducted experiments on both our dataset and the validation set of the VISOR dataset. Note that we used the VISOR validation set because, at the time of writing, the test set has not been released publicly. Finally, considering that active object categories weren’t labelled in VISOR, to ensure fairness in our comparison, we have employed our approach with an agnostic setting, hence without considering object categories at inference time. In all configurations, our approach outperformed the VISOR baseline by a significant margin. Specifically, for the test set EgoISM-HOI-Real we observed +9.68%, +6.59%, +4.53% when EgoISM-HOI-Synth, EgoISM-HOI-Real, EgoISM-HOI are used as training sets respectively. Furthermore, considering the validation set of the VISOR dataset, our approach surpassed the VISOR baseline by +2.21%. These results further confirm the effectiveness of the proposed method.